OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 3x - 1= (8x - 2) = 4 - 2 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3 Which step was the error made and why? (1 point) Question 19 options: 1) Step 1, because incorrect signs have been used 2) Step 2, because the terms have been added incorrectly 3) Step 3, because numbers have been added to both sides of the equation 4) No error was made

OpenStudy (anonymous):

can you please rewrite the question?

OpenStudy (anonymous):

sure @FriedRice

OpenStudy (anonymous):

im not sure if the equation 3x - 1= (8x - 2) = 4 is correct

OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 3x - 1= (8x - 2) = 4 - 2 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3 What Step was the Error made in? 1) Step 1, because incorrect signs have been used 2) Step 2, because the terms have been added incorrectly 3) Step 3, because numbers have been added to both sides of the equation 4) No error was made

OpenStudy (anonymous):

@FriedRice its supposed to be 1/2 like .50

OpenStudy (anonymous):

oh but is there suppose to be 2 equal signs 3x - 1/2 = (8x - 2) = 4

OpenStudy (anonymous):

or one of them has to be a plus sign?

OpenStudy (anonymous):

This is what it looks like @FriedRice

OpenStudy (anonymous):

ohh ok now it makes more sense thanks

OpenStudy (anonymous):

No Problem :]

OpenStudy (anonymous):

\[3x -\frac{ 1 }{ 2 }(8x - 2) = 4\]

OpenStudy (anonymous):

first you must distribute \[3x - 4x + 1 = 4\]

OpenStudy (anonymous):

then you must subtract or add common terms \[-x + 1 = 4\]

OpenStudy (anonymous):

or x = -3

OpenStudy (anonymous):

so from the statement what do you think is wrong or its all correct?

OpenStudy (anonymous):

So the sign was wrong it was supposed to be negative?

OpenStudy (anonymous):

yes so the from the explanation from the question step 2 made the mistake

OpenStudy (anonymous):

I don;t understand what is happening in step 3 though because a number comes in that wasnt there before.

OpenStudy (anonymous):

well its just subtracting both sides

OpenStudy (anonymous):

its not adding a number

OpenStudy (anonymous):

its just subtracting the original number and then moving the number to the other side

OpenStudy (anonymous):

Ohh I see because im usually used to it being subtracted vertically.